Two heterozygous red-eyed flies are crossed. The results are shown below. Determine the chi-square value and round to the nearest hundreth. Expected ratio is 3:1

Red eyes - 134

White eyes - 66
Answer #1

Let R be the allele for red-eye and r for white eye

The expected ratio of red to the white eye is 3:1.

or 3/4 red-eye flies and 1/4 white eye flies

Total progenies = 134+ 66 = 200

Expected red-eye flies - 3/4*200 = 150

Expected white-eye flies - 1/4*200 = 50.

Observed Expected O-E (O-E)2 (O-E)2/E
Red 134 150 -16 256 1.706667
White 66 50 16 256 5.12
Total 200 6.826667

So the Chi-square value is 6.826.

Please write to me if you need more help.